# Combinatorial Optimization using Comparison Oracles

### v1

@@ 3.2 Efficient Implementation of Global Subspace Learning
Note that our matrix entries are of the form $\mathbbm{1}_{S}-\mathbbm{1}_{T}$ where $S,T\in\mathcal{F}$, and hence they lie in $\{-1,0,1\}$. Since $m\leq n$, Lemma˜3.3 implies that the resulting integer basis vectors have entries bounded by $2^{\mathrm{poly}(n)}$. Let $W\in\mathbb{Z}^{s\times n}$ denote the matrix whose rows form this basis spanning the orthogonal lattice $\mathcal{W}$. The next implication is immediate:
Observation 3.4.
@@ 3.2.1 Efficient Separation via Dynamic Programming
as the set of vector values obtained by adding a subset of the first $i$ columns of $W$. We can compute $\mathcal{V}_{i}=\mathcal{V}_{i-1}\cup\{w_{i}+v:v\in\mathcal{V}_{i-1}\}$ starting from $V_{0}=\emptyset$ and increasing $i$ stopping when $|\mathcal{V}_{i}|>\ell$. This implies that there exists a value $v\in V_{i}$ that is not in $Z$ as $|Z|=\ell$.
The set $S$ such that $w(S)=v$ can be constructed by tracking the computation of $V_{i}$.
@@ 3.2.2 Efficient Separation for Linear Matroids via Algebraic Techniques
Given a linear matroid $\mathcal{M}=(U,\mathcal{I})$, non-negative integers $c_{1},c_{2},\dots,c_{n}$ for elements in $U$, and a set of values $Z=\{z_{1},\dots,z_{\ell}\}$, we can identify a basis $S\in\mathcal{F}$ of $\mathcal{M}$ such that $c(S)\notin Z$ (or decide that no such basis exists) in $\operatorname{poly}(n,\ell,\log M)$ time, where $M=\max_{i\in[n],j\in[\ell]}\{c_{i},z_{j}\}$.
@@
Warmup Proof of Lemma˜3.7.
@@
Furthermore, the associated a basis can be found using self-reducibility for matroids. For each element $j\in U$, we compute the “edge-deletion” polynomial $P_{j}(x)=\det(\sum_{i\neq j}x^{c_{i}}v_{i}v_{i}^{T})$: if $c$ appears as an exponent in $P_{j}(x)$, a basis with cost $c$ exists without $j$. We delete $j$ and recurse on the smaller problem. On the other hand, if $c$ does not appear in $P_{j}(x)$, then $j$ must be in every basis $S$ with $c(S)=c$. We select $j$ and recurse, searching for a basis with cost $c-c_{j}$ in the problem on $U/\{j\}$ (i.e., using an “edge-contraction” polynomial $P_{/j}(x)$) ∎
When the bound $M$ is large, the $\operatorname{poly}(n,M)$ time bound is infeasible, as the degree of $P(x)$ is too high for efficient interpolation. The full proof, which appears in Appendix˜B, avoids this problem by remaindering; i.e., working with the polynomial $P(x)\pmod{(x^{p}-1)}$ for a suitable prime $p\in\operatorname{poly}(n,\ell,\log M)$. This prime is chosen so that the residues modulo $p$ of all values in $Z$ and at least one basis cost $c\notin Z$ (if such a basis exists) are distinct. This allows for efficient interpolation and reconstruction within the required time bounds. Now we can use Lemma˜3.7 to prove Theorem˜1.4.
See 1.4
@@
From Algorithm Section˜3.1.1, the stated query complexity follows directly from Theorem˜3.2. It remains to bound the total running time required to implement the separation steps.
The maintenance of buckets, affine subspaces, and the corresponding updates can each be carried out in $\mathrm{poly}(n,B)$ time overall. By Lemma˜3.1, the number of separation steps is bounded by $O(nB)$. Each separation step invokes the procedures in Lemma˜3.6 and Lemma˜3.7, which take $\mathrm{poly}(n,\log M)$ and $\mathrm{poly}(n,\ell,\log M)$ time, respectively. Summing over all separation steps gives a total time of $\mathrm{poly}(n,B)$, using the facts that $\ell=O(nB)$ and $M=2^{\mathrm{poly}(n)}$, the latter following from Lemma˜3.3. ∎
@@ 3.3 Other Applications
In this section, we present several applications of Theorem˜1.4 and Theorem˜3.2 that yield low-query and efficient algorithms for well-studied problems such as $k$-SUM, SUBSET-SUM, and sorting sumsets $A+B$ (see [KLM19]). When the elements of the ground set take integer values in the range [B,B][-B,B], our algorithms achieve better query complexity while also being the first efficient algorithms in this setting.
@@ 3.3.1 Low-query and efficient comparison decision trees
    Using Theorem˜1.4 with $C=kB$, we obtain an algorithm with query complexity $O((n+kB)\log(kB))$ and running time $\operatorname{poly}(n,B)$, which improves both measures when $B=o(n)$. Our algorithm first sorts all feasible $k$-sets by their weights using only comparison queries between solutions, and then performs a binary search using comparisons with 0 to check whether a subset of size $k$ has total weight zero.
@@
    Using Algorithm Section˜3.1.1 with the efficient separation step stated in Section˜3.2.1, we obtain an algorithm with query complexity $O(nB\log(nB))$ and running time $\operatorname{poly}(n,B)$, which again improves both parameters when $B=o(n)$. As before, the algorithm first sorts all subset weights using comparison queries and then performs a binary search using comparisons with $t$ to determine whether a subset with total weight $t$ exists.
@@
    To fit this problem into our framework, we set $U=A\cup B$, and let $\mathcal{F}$ consist of the bases of a partition matroid over $U$ with parts $A$ and $B$, each having capacity one. Then comparing $w^{*}(a_{1})+w^{*}(b_{1})$ with $w^{*}(a_{2})+w^{*}(b_{2})$ corresponds exactly to comparing the weights of two feasible sets $\{a_{1},b_{1}\}$ and $\{a_{2},b_{2}\}$. Since partition matroids are representable over the reals, and the element values lie in [B,B][-B,B], we can apply Theorem˜1.4 with $C=2B$ to obtain an algorithm with query complexity $O((n+B)\log B)$ and running time $\operatorname{poly}(n,B)$. This improves on previous bounds when $B=o(n)$ and uses only valid pairwise comparisons among elements of $A+B$.
@@ 4 Minimum Cut using Cut Comparisons
We now turn to specific combinatorial problems, and give efficient algorithms using their structural properties. We begin with the min-cut problem on graphs and prove Theorems˜1.5 and 1.6, which we restate for convenience.
See 1.5
@@ 4.1 Graph Reconstruction
With the power of Theorem˜4.1 behind us, the graph recovery problem is easily solved. Indeed, using Theorem˜4.1(i), we can discover the neighborhood of node $u\in V$ using $n-1+O(\log n)$ queries and $\widetilde{O}(n)$ time; summing over all nodes gives us the graph $G$ in at most $O(n^{2})$ queries and $\widetilde{O}(n^{2})$ time.
To get the better bound of $O(m\log n)$ for sparse graphs, we use Theorem˜4.1(ii) to prove:
@@
Start with $T_{0}:=T$ and extract a neighbor $v_{i}\in N(u)\cap T_{i}$ (if any) from $\partial(u,T_{i})$ using Theorem˜4.1(ii); set $T_{i+1}:=T_{i}-v_{i}$, and repeat. Each step requires $O(\log n)$ queries and $\widetilde{O}(n)$ time. Repeating this for $r=\min(k,\partial(u,T))$ neighbors gives a total of $O(r\log n)$ queries and $\widetilde{O}(nr)$ time. A more efficient implementation, which amortizes the search and is detailed in Appendix˜C, achieves the claimed $\widetilde{O}(n)$ time. ∎
@@
Proof of Theorem˜1.6.
When $n\geq 4$, using Lemma˜4.3, we can discover all the neighbors of $u$ using $|\partial(u)|\cdot O(\log n)$ queries and $\widetilde{O}(n)$ time. Summing over all $u$ gives a total of at most $O((m+n)\log n)$ queries and $\widetilde{O}(n^{2})$ time. Hence, we start by running this procedure until the number of edges discovered is at least $n^{2}/\log n$. At this point, we switch to the algorithm using Theorem˜4.1(i) to discover the remaining edges using at most $O(n^{2})$ queries. This guarantees that we use at most $O(\min((m+n)\log n,n^{2}))$ queries and $\widetilde{O}(n^{2})$ time.
@@ 4.2 Finding Minimum Cuts
We now turn to finding minimum cuts. We start off by observing that when $n\leq 4$, we can bruteforce the minimum cut and otherwise, use Theorem˜1.6 to optimize over min-cuts: we reconstruct the graph using $O(\min((m+n)\log n,n^{2}))$ queries and $\widetilde{O}(n^{2})$ time; we can then optimize over it in $\widetilde{O}(n^{2})$ time. However, the number of cut queries incurred this way is quadratic and not near-linear, as promised in Theorem˜1.5.
However, we can use ideas from [RSW18] to do better. Suppose $G^{\prime}$ is a contracted graph (i.e., where some nodes in $G$ have been contracted into each other), and $G^{\prime}(p)$ is the “edge-percolation” graph obtained by sub-sampling each edge of $G^{\prime}$ independently with probability $p$. Distilling the results of [RSW18] shows that if, for any contraction $G^{\prime}$ of $G$, we can obtain a sample from $G^{\prime}(p)$ using $\alpha\cdot p\cdot|E(G^{\prime})|$ queries, then we can compute the min-cut of $G$ using $\alpha\cdot\widetilde{O}(n)$ queries and in $\widetilde{O}(n^{2})$ expected runtime. (A proof appears in Section˜C.2.)
@@
To sample $\partial_{G^{\prime}}(u,q)$, let $T$ be a random subset of $U-V_{u}$ with each element sampled independently with probability $q$. Reveal all the edges of $u$ going into $T$ using Lemma˜4.3. This takes $|\partial(u,T)|\cdot O(\log n)$ queries which is $q\cdot|\partial(u,U-V_{u})|\cdot O(\log n)$ in expectation and $\widetilde{O}(n)$ time. Summing over all $u\in U$ gives a total of $|E^{\prime}|\cdot O(p\log n)+\widetilde{O}(n)$ queries and $\widetilde{O}(n^{2})$ time in expectation. An edge $\{u,v\}$ for $v\in U\backslash V_{u}$ is sampled iff $v\in T$. This implies that each edge is sampled independently with probability $q$. Simply replacing $U$ with $S$ gives the extension to induced subgraphs. ∎
Combining Lemma˜4.4 with the abstraction of [RSW18] discussed above completes our algorithm for min-cuts (Theorem˜1.5).
@@ 4.3 Proof of the Structure Primitives Theorem
Finally, we turn to the proof of the Structure Primitives result in Theorem˜4.1. We will provide a proof sketch here for the query complexity bounds; see Appendix˜C for the full proof including the running time bounds.
Proof sketch of Theorem˜4.1.
@@
How do we find a median set for $u$? Order all the vertices other than $u$ (say this is $v_{1},v_{2},\ldots,v_{n-1}$, define $S_{i}$ to be the first $i$ elements. If $u$ has at least one edge incident to it, the query $\partial(S_{0}+u)-\partial(S_{0})$ has a positive sign, whereas $\partial(S_{n-1}+u)-\partial(S_{n-1})$ has a negative sign (note that we use $S_{0}=\emptyset$ and $S_{n-1}+u=U$ here for the sake of exposition, but such trivial cuts are not allowed in the actual comparison model. The full proof in Appendix˜C handles this correctly and with full rigor). So we can perform binary search to find a point where the sign changes, which can be used to find a median set in $\log_{2}n$ comparisons. (In fact, we can find two disjoint median sets—a prefix and a suffix of this ordering—and use the two to find all the neighbors in some $k$-sized set with $k+O(\log n)$ queries.)
@@ 4.4 Sparsifiers for Heavy Subgraph Problems
Because edges are sampled i.i.d. from the Bernoulli $\text{Ber}(p_{t})$ distribution at each step $t$, the symmetry implies that the subsample of edges is a uniform sample of $k$ edges. Using Lemma˜4.4, each sample $G(p_{t})$ takes $p_{t}\cdot\widetilde{O}(m)+\widetilde{O}(n)$ queries and $\widetilde{O}(n^{2})$ time in expectation. Summing this over $0\leq t\leq r$ gives an upper bound of $2p_{r}\cdot\widetilde{O}(m)+\widetilde{O}(nr)$ queries and $\widetilde{O}(n^{2}r)$ time. Using the fact that the probability that $G(p)$ has less than $k$ edges for $p=2k/m$, is at most $e^{-k/4}$, we can conclude that $p_{r}\leq 4k/m$ with probability at least $1-e^{-k/4}$. So the expected number of queries is upper bounded by $\widetilde{O}(nr+k)=\widetilde{O}(n+k)$ and the running time by $\widetilde{O}(n^{2}r)=\widetilde{O}(n^{2})$ using the fact that $r\leq O(\log n)$. ∎
Using Lemma˜4.5 with $k=\widetilde{O}(n/\varepsilon^{2})$, we can obtain a random subgraph $H$ of the unknown graph $G$ to get such a sample of edges using $\widetilde{O}(n/\varepsilon^{2})$ queries, and hence apply existing approximations for all heavy subgraph problems.
The results in this section do not generalize to the weighted case when the edges of the graph have unknown non-negative weights. We address this issue by giving some preliminary results under additional structural assumptions and slightly stronger oracle models—see Section˜C.3—but the general question remains wide open.
See Section˜C.4 for some illustrative examples which show some separation between the capabilities of the cut comparison and cut value query oracles.

### v2

@@ 4.2 Efficient Implementation of Global Subspace Learning
Note that our matrix entries are of the form $\mathbbm{1}_{S}-\mathbbm{1}_{T}$ where $S,T\in\mathcal{F}$, and hence they lie in $\{-1,0,1\}$. Since $m\leq n$, Lemma˜4.3 implies that the resulting integer basis vectors have entries bounded by $2^{\mathrm{poly}(n)}$. Let $W\in\mathbb{Z}^{s\times n}$ denote the matrix whose rows form this basis spanning the orthogonal lattice $\mathcal{W}$. The next implication is immediate:
Observation 4.4.
@@ 4.2.1 Efficient Separation via Dynamic Programming
as the set of vector values obtained by adding a subset of the first $i$ columns of $W$. We can compute $\mathcal{V}_{i}=\mathcal{V}_{i-1}\cup\{w_{i}+v:v\in\mathcal{V}_{i-1}\}$ starting from $\mathcal{V}_{0}=\emptyset$ and increasing $i$ stopping when $|\mathcal{V}_{i}|>\ell$. This implies that there exists a value $v\in\mathcal{V}_{i}$ that is not in $Z$ as $|Z|=\ell$.
The set $S$ such that $w(S)=v$ can be constructed by tracking the computation of $\mathcal{V}_{i}$.
@@ 4.2.2 Efficient Separation for Linear Matroids via Algebraic Techniques
Given a linear matroid $\mathcal{M}=(U,\mathcal{I})$, non-negative integers $c_{1},c_{2},\dots,c_{n}$ for elements in $U$, and a set of non-negative values $Z=\{z_{1},\dots,z_{\ell}\}$, we can identify a basis $S\in\mathcal{F}$ of $\mathcal{M}$ such that $c(S)\notin Z$ (or decide that no such basis exists) in $\operatorname{poly}(n,\ell,\log M)$ time, where $M=\max_{i\in[n],j\in[\ell]}\{c_{i},z_{j}\}$.
@@
Warmup Proof of Lemma˜4.7.
@@
Furthermore, the associated basis can be found using self-reducibility for matroids. For each element $j\in U$, we compute the “edge-deletion” polynomial $P_{j}(x)=\det(\sum_{i\neq j}x^{c_{i}}v_{i}v_{i}^{T})$: if $c$ appears as an exponent in $P_{j}(x)$, a basis with cost $c$ exists without $j$. We delete $j$ and recurse on the smaller problem. On the other hand, if $c$ does not appear in $P_{j}(x)$, then $j$ must be in every basis $S$ with $c(S)=c$. We select $j$ and recurse, searching for a basis with cost $c-c_{j}$ in the problem on $U/\{j\}$ (i.e., using an “edge-contraction” polynomial $P_{/j}(x)$) ∎
When the bound $M$ is large, the $\operatorname{poly}(n,M)$ time bound is infeasible, as the degree of $P(x)$ is too high for efficient interpolation. The full proof, which appears in Appendix˜C, avoids this problem by remaindering; i.e., working with the polynomial $P(x)\pmod{(x^{p}-1)}$ for a suitable prime $p\in\operatorname{poly}(n,\ell,\log M)$. This prime is chosen so that the residues modulo $p$ of all values in $Z$ and at least one basis cost $c\notin Z$ (if such a basis exists) are distinct. This allows for efficient interpolation and reconstruction within the required time bounds. Now we can use Lemma˜4.7 to prove Theorem˜1.5.
See 1.5
@@
From Algorithm Section˜4.1.1, the stated query complexity follows directly from Theorem˜4.2. It remains to bound the total running time required to implement the separation steps.
The maintenance of buckets, affine subspaces, and the corresponding updates can each be carried out in $\mathrm{poly}(n,B)$ time overall. By Lemma˜4.1, the number of separation steps is bounded by $O(nB)$. Each separation step invokes the procedures in Lemma˜4.6 and Lemma˜4.7, which take $\mathrm{poly}(n,\log M)$ and $\mathrm{poly}(n,\ell,\log M)$ time, respectively. Summing over all separation steps gives a total time of $\mathrm{poly}(n,B)$, using the facts that $\ell=O(nB)$ and $M=2^{\mathrm{poly}(n)}$, the latter following from Lemma˜4.3. ∎
@@ 4.3 Other Applications
In this section, we present several applications of Theorem˜1.5 and Theorem˜4.2 that yield low-query and efficient algorithms for well-studied problems such as $k$-SUM, SUBSET-SUM, and sorting sumsets $A+B$ (see [KLM19]). When the elements of the ground set take integer values in the range [B,B][-B,B], our algorithms achieve better query complexity while also being the first efficient algorithms in this setting.
@@ 4.3.1 Low-query and efficient comparison decision trees
    Using Theorem˜1.5 with $C=kB$, we obtain an algorithm with query complexity $O((n+kB)\log(kB))$ and running time $\operatorname{poly}(n,B)$, which improves both measures when $B=o(n)$. Our algorithm first sorts all feasible $k$-sets by their weights using only comparison queries between solutions, and then performs a binary search using comparisons with 0 to check whether a subset of size $k$ has total weight zero.
@@
    Using Algorithm Section˜4.1.1 with the efficient separation step stated in Section˜4.2.1, we obtain an algorithm with query complexity $O(nB\log(nB))$ and running time $\operatorname{poly}(n,B)$, which again improves both parameters when $B=o(n)$. As before, the algorithm first sorts all subset weights using comparison queries and then performs a binary search using comparisons with $t$ to determine whether a subset with total weight $t$ exists.
@@
    To fit this problem into our framework, we set $U=A\cup B$, and let $\mathcal{F}$ consist of the bases of a partition matroid over $U$ with parts $A$ and $B$, each having capacity one. Then comparing $w^{*}(a_{1})+w^{*}(b_{1})$ with $w^{*}(a_{2})+w^{*}(b_{2})$ corresponds exactly to comparing the weights of two feasible sets $\{a_{1},b_{1}\}$ and $\{a_{2},b_{2}\}$. Since partition matroids are representable over the reals, and the element values lie in [B,B][-B,B], we can apply Theorem˜1.5 with $C=2B$ to obtain an algorithm with query complexity $O((n+B)\log B)$ and running time $\operatorname{poly}(n,B)$. This improves on previous bounds when $B=o(n)$ and uses only valid pairwise comparisons among elements of $A+B$.
@@ 5 Minimum Cut using Cut Comparisons
We now turn to specific combinatorial problems, and give efficient algorithms using their structural properties. We begin with the min-cut problem on graphs and prove Theorems˜1.6 and 1.7, which we restate for convenience.
@@ 5.1 Graph Reconstruction
With the power of Theorem˜5.1 behind us, the graph recovery problem is easily solved. Indeed, using Theorem˜5.1(i), we can discover the neighborhood of node $u\in V$ using $n-1+O(\log n)$ queries and $\widetilde{O}(n)$ time; summing over all nodes gives us the graph $G$ in at most $O(n^{2})$ queries and $\widetilde{O}(n^{2})$ time.
To get the better bound of $O(m\log n)$ for sparse graphs, we use Theorem˜5.1(ii) to prove:
@@
Start with $T_{0}:=T$ and extract a neighbor $v_{i}\in N(u)\cap T_{i}$ (if any) from $\partial(u,T_{i})$ using Theorem˜5.1(ii); set $T_{i+1}:=T_{i}-v_{i}$, and repeat. Each step requires $O(\log n)$ queries and $\widetilde{O}(n)$ time. Repeating this for $r=\min(k,\partial(u,T))$ neighbors gives a total of $O(r\log n)$ queries and $\widetilde{O}(nr)$ time. A more efficient implementation, which amortizes the search and is detailed in Appendix˜D, achieves the claimed $\widetilde{O}(n)$ time. ∎
@@
Proof of Theorem˜1.7.
When $n\geq 4$, using Lemma˜5.3, we can discover all the neighbors of $u$ using $|\partial(u)|\cdot O(\log n)$ queries and $\widetilde{O}(n)$ time. Summing over all $u$ gives a total of at most $O((m+n)\log n)$ queries and $\widetilde{O}(n^{2})$ time. Hence, we start by running this procedure until the number of edges discovered is at least $n^{2}/\log n$. At this point, we switch to the algorithm using Theorem˜5.1(i) to discover the remaining edges using at most $O(n^{2})$ queries. This guarantees that we use at most $O(\min((m+n)\log n,n^{2}))$ queries and $\widetilde{O}(n^{2})$ time.
@@ 5.2 Finding Minimum Cuts
We now turn to finding minimum cuts. We start off by observing that when $n\leq 4$, we can bruteforce the minimum cut and otherwise, use Theorem˜1.7 to optimize over min-cuts: we reconstruct the graph using $O(\min((m+n)\log n,n^{2}))$ queries and $\widetilde{O}(n^{2})$ time; we can then optimize over it in $\widetilde{O}(n^{2})$ time. However, the number of cut queries incurred this way is quadratic and not near-linear, as promised in Theorem˜1.6.
However, we can use ideas from [RSW18] to do better. Suppose $G^{\prime}$ is a contracted graph (i.e., where some nodes in $G$ have been contracted into each other), and $G^{\prime}(p)$ is the “edge-percolation” graph obtained by sub-sampling each edge of $G^{\prime}$ independently with probability $p$. Distilling the results of [RSW18] shows that if, for any contraction $G^{\prime}$ of $G$, we can obtain a sample from $G^{\prime}(p)$ using $\alpha\cdot p\cdot|E(G^{\prime})|$ queries, then we can compute the min-cut of $G$ using $\alpha\cdot\widetilde{O}(n)$ queries and in $\widetilde{O}(n^{2})$ expected runtime. (A proof appears in Section˜D.2.)
@@
To sample $\partial_{G^{\prime}}(u,q)$, let $T$ be a random subset of $U-V_{u}$ with each element sampled independently with probability $q$. Reveal all the edges of $u$ going into $T$ using Lemma˜5.3. This takes $|\partial(u,T)|\cdot O(\log n)$ queries which is $q\cdot|\partial(u,U-V_{u})|\cdot O(\log n)$ in expectation and $\widetilde{O}(n)$ time. Summing over all $u\in U$ gives a total of $|E^{\prime}|\cdot O(p\log n)+\widetilde{O}(n)$ queries and $\widetilde{O}(n^{2})$ time in expectation. An edge $\{u,v\}$ for $v\in U\backslash V_{u}$ is sampled iff $v\in T$. This implies that each edge is sampled independently with probability $q$. Simply replacing $U$ with $S$ gives the extension to induced subgraphs. ∎
Combining Lemma˜5.4 with the abstraction of [RSW18] discussed above completes our algorithm for min-cuts (Theorem˜1.6).
@@ 5.3 Proof of the Structure Primitives Theorem
Finally, we turn to the proof of the Structure Primitives result in Theorem˜5.1. We will provide a proof sketch here for the query complexity bounds; see Appendix˜D for the full proof including the running time bounds.
Proof sketch of Theorem˜5.1.
@@
How do we find a median set for $u$? Order all the vertices other than $u$ (say this is $v_{1},v_{2},\ldots,v_{n-1}$, define $S_{i}$ to be the first $i$ elements. If $u$ has at least one edge incident to it, the query $\partial(S_{0}+u)-\partial(S_{0})$ has a positive sign, whereas $\partial(S_{n-1}+u)-\partial(S_{n-1})$ has a negative sign (note that we use $S_{0}=\emptyset$ and $S_{n-1}+u=U$ here for the sake of exposition, but such trivial cuts are not allowed in the actual comparison model. The full proof in Appendix˜D handles this correctly and with full rigor). So we can perform binary search to find a point where the sign changes, which can be used to find a median set in $\log_{2}n$ comparisons. (In fact, we can find two disjoint median sets—a prefix and a suffix of this ordering—and use the two to find all the neighbors in some $k$-sized set with $k+O(\log n)$ queries.)
@@ 5.4 Sparsifiers for Heavy Subgraph Problems
Because edges are sampled i.i.d. from the Bernoulli $\text{Ber}(p_{t})$ distribution at each step $t$, the symmetry implies that the subsample of edges is a uniform sample of $k$ edges. Using Lemma˜5.4, each sample $G(p_{t})$ takes $p_{t}\cdot\widetilde{O}(m)+\widetilde{O}(n)$ queries and $\widetilde{O}(n^{2})$ time in expectation. Summing this over $0\leq t\leq r$ gives an upper bound of $2p_{r}\cdot\widetilde{O}(m)+\widetilde{O}(nr)$ queries and $\widetilde{O}(n^{2}r)$ time. Using the fact that the probability that $G(p)$ has less than $k$ edges for $p=2k/m$, is at most $e^{-k/4}$, we can conclude that $p_{r}\leq 4k/m$ with probability at least $1-e^{-k/4}$. So the expected number of queries is upper bounded by $\widetilde{O}(nr+k)=\widetilde{O}(n+k)$ and the running time by $\widetilde{O}(n^{2}r)=\widetilde{O}(n^{2})$ using the fact that $r\leq O(\log n)$. ∎
Using Lemma˜5.5 with $k=\widetilde{O}(n/\varepsilon^{2})$, we can obtain a random subgraph $H$ of the unknown graph $G$ to get such a sample of edges using $\widetilde{O}(n/\varepsilon^{2})$ queries, and hence apply existing approximations for all heavy subgraph problems.
The results in this section do not generalize to the weighted case when the edges of the graph have unknown non-negative weights. We address this issue by giving some preliminary results under additional structural assumptions and slightly stronger oracle models—see Section˜D.3—but the general question remains wide open.
See Section˜D.4 for some illustrative examples which show some separation between the capabilities of the cut comparison and cut value query oracles.